what is the formula for  finding angle  of inner &  outer  of a any polygon

Grade:8

4 Answers

Arun
5 years ago
Dear student
 
Inner angle of a regular polygon = (n-2)* 180°/n
You can check yoirself by putting value of n = 1,2,3,.....etc
 
Regards
Arun
chirag
5 years ago
Inner angle of a regular polygon = (n-2)* 180°/n
You can check yoirself by putting value of n = 1,2,3,.....etc
Regards
chirag
 
Aishwarya
5 years ago
For sum of interior angles of polygon with side n = (2n - 4 )× 90
As per sum of external, it's always 360° 
If you've number of sides and you've find measure of each interior angle = 360/n 
Hope that helps! 
Sayantan Garai
5 years ago
Inner angle of a polygon= (n-2)×90°
n is the no. of sides.